OpenStudy (anonymous):

George and Karmen went on a bike trip. They took a bus to their starting point, and then biked the rest. They traveled 350 kilometers in total, and they biked 80 kilometers more than they were bussed. Let x=kilometers traveled by bike y=kilometers traveled by bus

OpenStudy (anonymous):

Find how many kilometers they traveled by bike

OpenStudy (anonymous):

@phi @kirbykirby @AnswerMyQuestions @zepdrix @hartnn @ganeshie8 @TuringTest

OpenStudy (kirbykirby):

You can set up as: x + y = 350.. the distance travelled is 350 in total y + 80 = x ... since the amount done by bike is 80 more than the bus

OpenStudy (anonymous):

Ok so what should i do to get x and y?

OpenStudy (kirbykirby):

You can sole each equation for x (or y if you wanted), then substitute what x is equal into the other equation... Like: x + a = y -x + -y = 2 Then you can do x = y - a and x = -y - 2 then equate both equations: y - a = -y -2

OpenStudy (kirbykirby):

can solve*

OpenStudy (anonymous):

Haha im sorry.. youre explaining it well its just im confused. Im not very good with equations.

OpenStudy (kirbykirby):

Ok don't worry :) One step at a time.. First, are you able to solve each equation for x (meaning, isolate your x variable so that only one x is on one side of the equals sign... like " x = ... " )

OpenStudy (anonymous):

Ummm it all depends on the equation. But on this one i wouldnt know how

OpenStudy (kirbykirby):

For the second equation, your x is already on one side y + 80 = x same as: x = y + 80

OpenStudy (kirbykirby):

For the 1st equation: x + y = 350 you can subtract y from both sides, this will get rid of the y on the right side, and leave x only

OpenStudy (anonymous):

ok

OpenStudy (kirbykirby):

what do you get

OpenStudy (anonymous):

x=350 right..

OpenStudy (kirbykirby):

not quite, you have to subtract y on both sides.. so you should have: x = 350 - y

OpenStudy (anonymous):

I have 3 min do you know what the answer is??

OpenStudy (anonymous):

THANK YOU SO MUCH